A circuit has in series an electromotive force given by E=40-sin35tV, a resistor of 10Ω, and an inductor of 0.4H . If the initial current is 2 , find the current at time t>2.


1
Expert's answer
2021-12-16T07:27:28-0500

"X_L = wL =35\cdot0.4= 14\ Ω"

"Z = R + jwL = 10 + j14\ Ω"

The magnitude of Z:

"Z = \sqrt{R^2 + (wL)^2} = \sqrt{10^2 +14^2} = 17.20\ Ω"

The phase angle of Z:

"\varphi = tan^{-1} ( wL/R) = tan^{-1} (14/10) = 54°=0.94" rad


current:

"I(t)=2sin(35t-0.94)"

"I(2)=2sin(70-0.94)=-0.11" A